40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 232 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  5654   Wed Oct 12 00:35:42 2011 kiwamuUpdateLSCTRY path realigned

The TRY (TRansmitted light from Y arm ) path was a bit realigned because there had been a small clipping.

This clipping was introducing offsets on the error signals of the C1ASS servo.

(Story)

During I was running the C1ASS servo on the Y arm I found every time after the auto-alignment is done there still remained a slight offset in the beam pointing,

I looked at the CCD camera which looks at the transmitted light and then introduced an intentional misalignment in ETMY in order to find an obvious clipping.

Indeed there was a clipping in horizontal direction. I checked through the optics on the Y end optical bench.

On the second mirror (beam splitter) the beam was on a very edge. So I steered the first steering mirror to fix it,

In addition to that an iris which is placed between the first and second mirror was also clipping the beam,

So I fully opened the aperture of the iris.

  5653   Tue Oct 11 21:23:51 2011 JenneUpdateLSCArm absl lengths

Quote:

[Katrin, Jenne]

We took the data for the new absolute length measurement of both arms, after the latest vent and move.  We will analyze soonly.  We had done a round of analysis,  but then Koji pointed out that our data wasn't so clean because the whitening filters were on (and saturated the ADC).  We now have the data (but not the analysis) for the better data with the WF off.

So our dirty-data preliminary number for the X arm is 37.73meters, which is 14cm different from our old length.  We were supposed to move by ~20cm, so....either this measurement is bad because the data sucked (which it did), or we are 6cm off.  Or both.

I'll do another analysis with the clean data for both arms later today/tomorrow.

After analyzing the cleaner data, I get the following:


Y_Length_long  =  37.757 meters

X_Length_long  =  37.772 meters

 

As stated in the wiki, the goal arm length was  L = 37.7974 m for each arm. 

So we're within 2cm for X, and within 4cm for Y.

According to Kiwamu's awesome tolerance calculation, we need to be within 2cm for each arm.  Given that we started out 20cm wrong for X and 25cm wrong for Y, we're a lot closer now, even though we aren't meeting our Yarm requirement yet.

Probably some Optickle action is in order, to see what these new lengths give us in terms of sideband phase and other stuff.

If you want more digits on my calculated numbers (which are probably meaningless, but I haven't done a careful error analysis), in my directory ...../users/jenne/Xarm and ..../users/jenne/Yarm run Xarm_find_peaks_and_length.m and Yarm_find_peaks_and_length.m  respectively.  These will output the lengths.

  5652   Tue Oct 11 19:11:25 2011 kiwamuUpdateLSCRe: BS actuator reponse at low frequency : measured

I think the precision due to the loop gain uncertainty is something like 0.1% at 0.1 Hz. It's not the issue.

The real issue was the loud motion of MICH, which degrades the coherence of the measurement.

Also last night I tried the fringe hopping technique and gave it up for several reasons.

 

(uncertainty due to the loop gain)

When MICH is locked, the signal at C1:LSC-MICH_OUT can be expressed in frequency domain by
 
     MICH_OUT = G / (1+G) * (1 / A) * X + G / (1+G) * (1 / H) * (1 / A) * S,                 [1]
 
where G is the open loop gain, A is the actuator response, H is the sensor transfer function (constant factor),
X is the natural (unsuppressed) motion of MICH and S is an excitation injected at C1:LSC-MICH_EXC.
When the natural motion of MICH X is smaller than the excited displacement S/H, dividing MICH_OUT by S gives
 
   [Transfer function] = S / MICH_OUT
                                 = (1+G) /G * H * A
 
At low frequency the open loop gain is always big, so that the transfer function can be approximated to
 
   [Transfer function] ~ H *A
 
This approximation is valid with a precision of 1/G.
In my case yesterday, the open loop gain at 0.1Hz was about 103 or more than that, so the uncertainty due to the loop gain was 0.1% or even less.
 

(Effect from the MICH motion)

In the equation [1], it is shown that the MICH motion X shows up together with the excitation signal.
Actually this MICH motion term was not completely negligible and eventually this term disturbs the measurement resulting in a low coherence.
In order to get a high coherence in the measurement, X should be smaller than the excited displacement S/H,
 
      X << S / H
 
This the reason why I had to inject a big excitation signal. Although the coherence around 1Hz turned out to be still low due to the loud natural motion in MICH.
The excitation was already close to 0.1 um level in terms of peak-to-peak displacement, and I wasn't able to increase it any more because the MICH signal would run into a nonlinear regime.
In the worst case I lost the lock due to a too much excitation.
 
 
(Fringe hopping technique)
 
Actually I tried and gave up this technique. That's why I did the in-loop measurement.
My feeling is that this technique is not suitable for the 40m.
What I tried was to flip the sign of the MICH control such that the fringe hops from the dark fringe to the neighbor bright fringe or vice versa.
Difference in the control signal (C1:LSC-MICH_OUT) was supposed to give us the amount of signal which drives the actuator by exactly quarter of the laser wave length.
However this technique turned out to be not good because
    (1) BS actuator is too strong
          => expected difference in the control signal is quite small.
          => \lambda / 4 / A ~ 12 counts, where A is the actuator DC response of about 2.2e-8 [m/counts].
   (2) MICH motion was too loud
         => I saw such a tiny 12 counts difference in the control signal, but once the hopping is done the control signal immediately fluctuated and it was really hard to precisely measure it.
         => It's simply because MICH was loud, and the actuator tried to suppress the motion and it resulted such an immediate signal fluctuation in the control signal

Quote from #5649

This seems like an error prone method for DC responses due to the loop gain uncertainty. Better may be to use the fringe hopping method (c.f. Luca Matone) or the fringe counting method

 

  5651   Tue Oct 11 17:32:05 2011 jamieHowToEnvironment40m google maps link

Here's another useful link:

http://maps.google.com/maps?q=34.13928,-118.123756

  5650   Tue Oct 11 15:19:17 2011 ranaHowToEnvironment40m map

The Kinemetrics dudes are going to visit us @ 1:45 tomorrow (Wednesday) to check out our stacks, seismos, etc.

40mLabMap.png40mLabMap.jpg

I put these maps here on the elog since people are always getting lost trying to find the lab.

  5649   Tue Oct 11 15:14:50 2011 ranaUpdateLSCBS actuator reponse at low frequency : measured

Quote:

The response of the BS actuator in a low frequency regime has been measured.

 

This seems like an error prone method for DC responses due to the loop gain uncertainty. Better may be to use the fringe hopping method (c.f. Luca Matone) or the fringe counting method

  5648   Tue Oct 11 03:35:16 2011 kiwamuUpdateLSCBS actuator reponse at low frequency : measured

The response of the BS actuator in a low frequency regime has been measured.

After the measurement I did a coarse fit to see if the low frequency data agree with the high frequency response which I have measured two weeks ago (#5583)
So far it shows a good agreement with the high frequency data (see the plot below). Tomorrow I will do a serious fitting.
Once the calibration of BS is done, the low frequency responses of ITMs, PRM and SRM will be done by simply exciting BS and comparing them (maybe at a couple of frequency points around 0.1Hz).
BSactuator.png

(Measurement)

 + With free swinging MICH, the sensor (AS55_Q) was calibrated into counts/m.

     => The peak-peak counts was about 110 counts. So the sensor response is about 6.5x108 counts/m

 + Locked Michelson with AS55_Q and the signal was fedback to BS.

 + Set the UGF high enough so that the open loop gain below 10 Hz is greater than 1.

 + With DDT's swept sine measurement, C1:LSC-MICH_EXC was excited with a big amplitude of 40 counts.

 + Took a transfer function from C1:LSC-MICH_OUT to C1:LSC-MICH_EXC.

 + Calibrated the transfer function into m/counts by dividing it with the sensor response.

Quote from #5641
One possible reason is that my actuator responses are not so accurate below 1Hz.
I will measure the DC response of all the actuators and it will completely determine the shapes of the actuator responses except for the region around the resonance.

  5647   Tue Oct 11 00:59:55 2011 SureshUpdateIOOPreliminary locking of WFS loops

[Kiwamu, Koji, Suresh]

After correcting several errors in the WFS loops, we turned them on today and saw them working!

A while back (last week actually) I noticed that the WFS1 and WFS2  QPD segments were numbered in a different order but that their input matrices did not reflect this change.  As result the WFS pitch and yaw definitions were pretty much mixed up.  However even after clearing this up the signals still showed significant amount of cross couplings. 

This problem was finally traced to the relative phase between I and Q channels of the WFS segments.  Koji suggested that I check the relative phase between all the segments to be sure.  I then repeated the procedure that Valera and I followed in our earlier elog # 5321 , and found that the phases indeed required to be adjusted.  The excitation of MCL was at 6Hz, 100mVpp, as before.   The WFS response after this was much improved i.e.  the pitch yaw cross couplings were not visible when we misalign the MC with sliders in MC_ALIGN.  And the magnitude of the response also increased since the signal was transferred from the Q to I channels.  The the phases were tweaked by hand till Q< 1% of I.  However when I repeated this measurement an hour later (I wanted to save the plots) I found that the phases had changed by a few percent! 

Koji noticed that the MC_REFL camera image showed significant intensity fluctuations and advised that we try a higher frequency and lower amplitude to avoid nonlinear effects in the WFS and in the MCL to PSL lock.  So we repeated the process at 20Hz and 20mVpp, introduced at the IN2 of the MC_Servo.  The fig below shows the level to which we reduced the signal in Q.

WFS1_IQphase20111010.pdf    WFS2_IQphase20111010.pdf

We then checked the relative phase between various quadrants by looking at the time series in dataviewer.  WFS2 Seg4 phase had to be flipped to bring it into phase with all the rest. 

WFS_IQ_RelativePhase.png

 

After this I tried to see the WFS response to moving the MC1 and MC3 with the sliders and determined the following relations:

Pitch WFS1 WFS2
MC1 + -
MC2 - -
MC3 + +

 

Yaw WFS1 WFS2
MC1 + +
MC2 - -
MC3 + -

 

Disregarding the MC2 for now and assuming arbitrary gains of 1 for all elements we inverted these matrices inserted them into the WFS_servo_outmatrix.  We then found that the with a sign flip on all elements the loops were stable.  In the servo filters we had turned on only the filter modules 3 and 4.  There was no low frequency boost.   We gradually increased gain till we saw a significant suppression of the error signal at low frequencies as shown below.  There was also an associated suppression of Intensity noise at REFL_DC after a single bounce from PRM.

 WFS_error_signal_Oct10.pdf        WFS_reduction_intensity_noise_Oct10.png

 

To see if the locks can actually realign the MC if it were manually misaligned, we turned the loops off and misaligned MC by moving MC3 pitch by 0.05 (slider position), and then turned on the loops.  The locks were reengaged successfully and the MC regained alignment as seen on the StripTool below:

WFS_recentering_Oct10.png

 

We can now proceed with the fine tuning the servo filters and understand the system better:

Q1:    Does the WFS (I to Q) phase drift rapidly?  How can we prevent it?

Q2:   How is that we do not see any bounce or roll resonances on the WFS error signals?

Q3:  How do we include the MC2 QPD into the WFS Servo?

I will proceed with determination of the actual transfer coefs between the MC DoF and the WFS sensors. 

 

 

  5646   Mon Oct 10 18:53:04 2011 KatrinUpdateGreen LockingMirrors whose angle of incidence is not 45

The angle of incidence of light is for some mirrors on the YARM end table different from 45° even though the mirrors are coated for 45°.

The mirrors below are useful if there are plans to replace these mirrors by properly coated ones.

 

Mirror
Angle of incidence (degree)
1st 1" mirror right after laser* 10
2nd 1" mirror right after laser 35
1st 2" steering mirror to vacuum system 15
2nd 2" steering mirror to vacuum system 28

 

* This is the new mirror as decribed on http://nodus.ligo.caltech.edu:8080/40m/5623

 

  5645   Mon Oct 10 16:32:18 2011 steveUpdateSUSUL sensor of ETMY is recovered

 I lost UL osem voltage this morning when I was checking the actual connection at rack ETMY

This after noon I disconnected  the 64 pins IDE connector from satelite amp at the rack, and the two 25 pins Dsubs at this juction board.

UL OSEM recovered after reconnecting these three connectors.

Atm3, bad connection.........noisy UL

 

Attachment 1: ETMY_UL.png
ETMY_UL.png
Attachment 2: ETMY_OSEM_UL.png
ETMY_OSEM_UL.png
Attachment 3: noisyETMY_UL.png
noisyETMY_UL.png
  5644   Mon Oct 10 15:41:56 2011 KojiUpdatePSLPMC aligned

[Koji Suresh]

The steering mirrors for PMC were aligned. The transmission went up from 0.779 to 0.852.

  5643   Mon Oct 10 13:52:04 2011 kiwamuUpdateLSCRE: First attempt to estimate mode matching efficiency using interferometer

Quote from #5640

"^2"s are missing in the second equation, but the calculation results seem correct.

PRX and PRY have different mode matching because of the Michelson asymmetry.
Are individually estimated mode matching indicates any sign of reasonable mode mismatch?
(The difference can be very small because the asymmetry is not so big.)

- Thank you for the correction. The missing square operation has been added correctly on the last entry (#5639).

- As for the individual MM efficiency,
   I was assuming that the MM solutions are the same for PRX, PRY and the real PRC, so I haven't carefully checked differences between those cavities.
   However as you mentioned the difference in those cavities can be tiny due to the small 3 cm Schnupp asymmetry.
   Anyway I will briefly check it to make me sure.
  5642   Mon Oct 10 12:14:00 2011 MirkoUpdateComputer Scripts / ProgramsIMC simulations

[Mirko, Kiwamu]

I tried to answer two questions regarding the IMC:

1. What is the coupling of fluctuations in the SB freq. to SB transmitted power?
2. What (if any) is the influence of the IMC on the AM?

I ran into some weird things regarding the corresponding optickle simulations:
1. There seems to be some artifact at the beginning of every simulation sweep.
2. The position of features depends on the parameters of the sweep.

I mailed Matt asking if he sees some error in the simulations

opticke_xaxis.png

Attachment 2: DC_power.png
DC_power.png
Attachment 3: DC_power_B.png
DC_power_B.png
Attachment 4: IMC_simulation.zip
  5641   Mon Oct 10 10:14:43 2011 ranaUpdateLSClength fluctuations in SRCL

 How does it make sense that the motion at 0.1 Hz of PRC is 10x larger than MICH?

EDIT by KI:

 That's actually the point which I was wondering at. One possible reason is that my actuator responses are not so accurate below 1Hz.
I will measure the DC response of all the actuators and it will completely determine the shapes of the actuator responses except for the region around the resonance.
In the process of producing the plot I was assuming that all the actuator response have a 1 Hz resonance with Q of 5.
However in reality this assumption is not true because the resonant frequency is different in each actuator.
  5640   Mon Oct 10 00:01:26 2011 KojiUpdateLSCFirst attempt to estimate mode matching efficiency using interferometer

"^2"s are missing in the second equation, but the calculation results seem correct.

PRX and PRY have different mode matching because of the Michelson asymmetry.
Are individually estimated mode matching indicates any sign of reasonable mode mismatch?
(The difference can be very small because the asymmetry is not so big.)

  5639   Sun Oct 9 17:13:46 2011 kiwamuUpdateLSCFirst attempt to estimate mode matching efficiency using interferometer

The efficiency of the mode matching (MM) to PRC (Power-Recycling Cavity) has been estimated by using the interferometer.

The estimated MM efficiency is about 74 % when losses in the cavity are assumed to be zero.

 

(Motivation)

 There had been an issue that the recycling gain didn't go to the designed high value of about 42  (#5541).
One of the possibilities is a low efficiency in the MM to PRC (also see #5541).
Although the MM efficiency had been measured using a beam scan ( see a summary on the wiki) a long time ago, it haven't been verified.
Therefore the MM has to be reviewed by using the real interferometer.

(Measurement)

 The concept of this measurement is observe the amount of the reflected light from a power-recycled cavity and estimate the MM efficiency based on the measured reflectivities.
 Since using the real PRC (consisting of BS, ITMs and PRM) could be a too complicated system for this measurement,
simpler cavities, namely Power-Recycled ITMX and ITMY (PRX and PRY), were used to examine the MM efficiency.
 The measurement goes in the following order :
    (1) Measurement of the amount of the single-bounce reflection from PRM with BS and ITMs misaligned.
    (2) Lock PRX or PRY to carrier resonance.
    (3) Alignment of PRX/Y to maximize the intracavity power. This time ASDC was used as a monitor of the intracavity power.
    (4) Measurement of the amount of the reflected light when the cavity is in resonance. The value in REFLDC was averaged in 100 sec.
     => done by tdsavg 100 C1:LSC-REFLDC_OUT
The same measurement was performed for both PRX and PRY.
 
 - locking parameters -
  Sensor = REFL11_I
  Whitening gain = 10 (30 dB)
  PRCL_GAIN = 2
  UGF ~ 200 Hz

(Analysis)

In order to estimate the relation between the MM efficiency vs. the reflected light, two models are considered:
   (1) simple model => no loss and no sidebands
   (2) sideband-included model => no loss but sidebands are taken into the account of the reflection.
 
(1) In the simple model the reflectivity Prefl / Pin is expressed by
         [Reflectivity]  = Prefl / Pin = Z * Rcav +  (1- Z) * Rprm
 
where Z is MM efficiency and Rprm is the reflectivity of PRM
and Rcav is the reflectivity of PRX/Y when it's resonance and it is defined by
         Rcav = | rprm - ritm t2BS|2 / |1 -rprm ritm t2BS |2
 
Tprm = 5.75% and Titm = 1.4 % are assumed in all the calculations.
In the first equation the first term represents the mode matched light and hence it couples with the cavity through Rcav.
The second term is the non-mode-matched light and because they are not interacting with the cavity they will be simply reflected by PRM through Rprm.
 
(2) In reality two phase-modulated light (11 MHz and 55 MHz) will behave differently from the carrier.
For example when the carrier is in resonance the sidebands will be anti-resonance against the cavity.
So that the amount of REFLDC will be slightly bigger because of the reflection of the sidebands.
 
      Prefl = Z * Rcav * Pc + Z * Ranti * Ps +  (1- Z) * Rprm * (Pc + Ps)
 
where Pc and Ps are the power in the carrier light and the sidebands respectively.
And Ranti is the reflectivity of the anti-resonance PRX/Y, which can be obtained by replacing the minus sign by the plus sign in the equation of Rcav shown above.
It is assumed that the sum of the carrier power and sidebands power is the incident power Pin = Pc + Ps.
The power in the carrier and the sidebands were estimated based on the OSA measurement (#5519), so that
          Pc / Pin = |J0(0.14)|2 * |J0(0.17)|2 =  0.976
          Ps / Pin = 2 * |J1(0.14)|2 + 2 * |J1(0.17)|2 =  0.024
 

(Results)

Here are the measured values in REFLDC

 -- Measurement 1 : PRX
    Single bounce from PRM = 4802.27 counts
       ==> the incident power = 5095.25 counts
    Reflected light from PRX = 4433.88 counts
      ==> Reflectivity = 0.8702
 
-- Measurement 2 : PRY
    Single bounce from PRM = 4833.05 counts
       ==> the incident power = 5127.05 counts
    Reflected light from PRX = 4444.48 counts
      ==> Reflectivity = 0.86672
 
On average the reflectivity of power-recycled ITM cavity was 0.868 with a standard deviation of  0.001744.
Actually the standard deviation estimated here is not fair because the measurement was done by only twice,
but my intention was that I wanted to see how the error can affect the estimation of the MM efficiency.
Here is a plot comparing the model curves and the measured values with 5 sigma error box (5 times of measured standard deviation).

mm_reflection.png

It is shown that the mode matching efficiency is 73.7 % when the sideband-included model is considered.
With the 5 sigma deviation it can go from 65% to 82% but it is still low and lower than the beam scan measurement ( see a summary on the wiki). 

Anyways the estimated MM efficiency with the sidebands effect included and without loss effect is

        MM efficiency = 73.7 +/- 1.7 % (1 sigma error)  or +/- 8.7 % (5 sigma error)

  5638   Sat Oct 8 04:41:07 2011 kiwamuUpdateLSClength fluctuations in SRCL

For a comparison, the length fluctuation of Signal-Recycled ITMX (SRX) and ITMY (SRY) have been measured.

Roughly speaking the length motion of SRX and SRY are as loud as that of PRCL.

Some details about the measurement and data analysis can be found in the past elog entry (#5582).

In the process of converting the raw spectra to the calibrated displacements the SRM actuator was assumed to have a resonance at 1Hz with Q = 5.

length_noise.png

(Notes on SRX/Y locking)

     Sensor = REFL11_I
     Actuator = SRM
     Demod. phase = 40 deg
     SRCL_GAIN = 20
     UGF = 100 - 200 Hz
     Resonant condition = Carrier resonance
     Whitening gain = 0 dB
     ASDC = 360 counts

Quote from #5582

The MICH and PRCL motions have been measured in some different configurations.

      + PRCL is always noisier than MICH.

  5637   Sat Oct 8 00:44:42 2011 kiwamuUpdateLSCcalibration of SRM actuator

The AC response of the SRM actuator has been calibrated.

 actuators.png
(Summary of the calibration results)
     BS = 2.190e-08 / f2     [m/counts]
     ITMX  = 4.913e-09 / f2   [m/counts]
     ITMY  = 4.832e-09 / f2   [m/counts]
     PRM   = 2.022e-08 / f2  [m/counts]
     SRM   = 2.477e-08 / f2  [m/counts]    ( NEW ! ) 
 
(Measurement)
The same technique as I reported some times ago (#4721) were used.
The Signal-Recycled ITMY was locked for measuring the actuator response.
Since the ITMY actuator had been already calibrated, first the sensor was calibrated into [counts/m] by exciting the ITMY actuator and then calibrated the SRM actuator with swept sine measurement.
 
 - - notes to myself
   SRCL GAIN = 2.2
   Sensor = REFL11_I
   Demod. phase = 40 deg
   Resonant condition = Carrier resonant
   Gain in WF = 0 dB

Quote from #5583
The AC responses of the BS, ITMs and PRM actuators have been calibrated.

 

  5636   Fri Oct 7 23:23:05 2011 kiwamuUpdateSUSSRM oplev was oscillating

The SRM oplevs were found to be oscillating because of a small phase margin.

I reduced the gains to the half of them. The peak which Steve found today maybe due to this oscillation.

Quote from #5630

The SRM bounce peak 18.33 Hz. Suresh helped me to retune filter through Foton, but we failed to remove it. 

  5635   Fri Oct 7 22:48:26 2011 SureshUpdateIOOMC2 Trans QPD spot size and incident power decreased

After centering the spot on the MC2, I started to adjust the spot position on MC_TRANS_QPD to center the beam on it.  I noticed that the spot size was about 3 to 4mm dia. because the 200mm lens was too close to the QPD.  I moved it back and decreased the spot size to about 1mm and the sensitivity to spot position increased.  However, Koji noted that the QPD sectors were near saturation, so I put in a ND=0.3 filter to reduce the incident power on the QPD.

At optimal alignment the current QPD_SUM is around 25k to 26k counts (factor of 2 down). Eventually the gain of the QPD ckts have to be reduced to prevent saturation, for the moment this is temporary fix.

The MC_TRANS_SUM trigger for MC autolocker is working fine no further change was required.

  5634   Fri Oct 7 22:41:05 2011 KojiConfigurationGeneralScript backup fixed

Script backup regularly runs on op340m by crontab,
but the true backup were not taken since Oct 16, 2010
as the backup program was looking at the old script directory.
/cvs/cds/script/backupScripts.pl was modified to look at the new script directory.

OLD COMMAND:

$command = "cd /cvs/cds/caltech; /usr/sbin/tar cfX - $EXCLUDE_LIST scripts | bzip2 > $ARCHIVEDIR/scripts_$curdate.tar.bz2";

NEW COMMAND:

$command = "cd /opt/rtcds/caltech/c1; /usr/sbin/tar cfX - $EXCLUDE_LIST scripts | bzip2 > $ARCHIVEDIR/scripts_$curdate.tar.bz2";

  5633   Fri Oct 7 22:31:53 2011 kiwamuUpdateSUSnoisy ULSEN on ETMY

Yesterday Koji was claiming that the rms monitor of ULSEN on ETMY didn't well go down.

Indeed something bad is happening on ULSEN of ETMY.

I guess it could be a loose connection.

ETMY_OSEMs.png

(The unit of Y axis in the plot is not true. Don't believe it !)

  5632   Fri Oct 7 19:06:46 2011 SureshUpdateIOOMC spot positions: checked and corrected.

Koji and Kiwamu had adjusted the MC beam axis slightly such that we can couple the MC output into the Y-arm without exceeding the current range of adjustability on PZT1.  This changed the centering of beam spots on MC mirrors.  I checked the mc-decentering make sure we have not made too big a compromise.  And since we can move MC2 spot position while maintaining the current positions on MC1 and MC3 decentering, we can atleast eliminate the A2L coupling on that mirror.  I used the scripts in $scripts$/MC/moveMC2/ to adjust the MC2 spot position.

Spot positions in mm (MC1,2,3 pit MC1,2,3 yaw) before adjustment:
    1.4674   -0.3548    1.0199   -1.5519    1.9834   -1.5971

After correcting MC2:

    1.4528    0.1431    0.9958   -1.2147    0.3823   -2.0163

After correcting MC1:

    1.3745    0.0669    0.8899   -1.5269    0.0296   -1.7314

 

The spot positions on MC1 and MC3 are very nearly (+/- 0.06 mm) same as before, while the MC2  decentering has been reduced close to zero.

A slight adjustment of the PZTs may be required to reset the beam pointing.

  5631   Fri Oct 7 17:35:26 2011 KatrinUpdateGreen LockingPower on green YARM table

After all realignment is finished, here are the powers at several positions:

 

DSC_3496_power.JPG

  5630   Fri Oct 7 14:04:48 2011 steveUpdateSUSHe/Ne intensity noise effect on oplevs

The SRM bounce peak 18.33 Hz. Suresh helped me to retune filter through Foton, but we failed to remove it.

ITMX_OPLEV_PERROR shows high coherence with oplev laser. This is our lousiest set up. I will work on it next week.

Generally speaking we can say that JDSU-Uniphase 1103P and 1125P laser intensity noise do not limit oplev servo performance.

However, the overall well being of filters, gain settings, beam sizes on qpds are poor.

 

Attachment 1: PRMoplevINTn.png
PRMoplevINTn.png
Attachment 2: BSoplevINTn.png
BSoplevINTn.png
Attachment 3: ITMXoplevINTn.png
ITMXoplevINTn.png
Attachment 4: ITMYoplevINTn.png
ITMYoplevINTn.png
Attachment 5: SRM_oplevINTn.png
SRM_oplevINTn.png
Attachment 6: ETMYoplevINTn3.png
ETMYoplevINTn3.png
Attachment 7: ETMXoplevINTn.png
ETMXoplevINTn.png
  5629   Fri Oct 7 11:53:47 2011 KojiUpdateLSCDRMI locked and some plans

- REFL165 PD to be fixed (shows constant high voltage at the DC out)

- Make POP22/110 PD

- Install AS11? or use it as POX11?

- Install POP55

  5628   Fri Oct 7 11:45:24 2011 KojiSummaryLSCPOY11 installed, 55MHz PD at POY removed

POY11 PD was installed last night. The lock of the Y arm was confirmed with the POY11I signal.

- The DC transimpedance was modified to be 1010V/A as the incident power is tiny.

- The demodulation phase of the roughly adjusted (148deg) to have PDH signal at the I-phase.
 
The comparison with AS55I signal exhibits that POY11I have ~150 times weaker signal with 45dB whitening.
 
(In total 25000 times weaker.)

On the way to make POY11 functioning, there were many fixes at the LSC rack...


Details:

- The PD interface cards (power supply for the RFPDs) were checked:
So far the two card at the right hand side were checked. 

Desipite the previous entry reported the issues on those boards, they did not show any problem yesterday.

One hypothetical possibility is the enabling switches that is controlled from the old slow epics targets.

- POY55 was removed

This 55MHz PD is supposed to be installed at POP.
The PD, an RF cable, an RF amp, the power supply of the RF amp were removed.

- POY11 was installed

The PD was placed where the 55MHz was placed.

The beam was aligned on the diode using the IR viewer and the digital multimeter.

The power supply cable and the RF cable for POY on the ITMY table were used.

There were an ND filter on the POY beam path. It was removed.


- On the LSC rack
The PD RF was connected to the patch panel at the top of the rack.

There were loose connectors on the patch panel. Some connectors were tightened on the panel.

I found that POY11 and POX11 had I&Q signal reversely connected to the whitening board.
   ==> These were fixed but
require the orthogonality test again for those channels.



The I phase output of the AS11 demod board had a broken connector. 

The onboard SMA has got disintegrated because of too much twist on the connector.
The board was once removed from the rack and the connector was fixed using a heat gun and soldering.

The DC signals were checked. POYDC was not correctly connected. POYDC were correctly connected to the POYDC channel.

- CDS
c1lsc was found with the RFM frozen.
The c1lsc machine was soft-rebooted after stopping all of the RT processes.

Once the RT processes came back, they were all burtrestored.

- PDH locking
Restored Y-arm. Locked it with AS55Q.
Ran ASS alignment for Y-arm.
100cnt 150Hz sinusoidal signal is applied to ETMY

Measured the PSD of AS55Q, POY11I, and POY11Q.
Adjusted the demod phase so that the excitation could be minimized in POY11Q.


  5627   Fri Oct 7 04:42:24 2011 kiwamuUpdateLSCDRMI locked and some plans

DRMI has been locked using the same RFPD selection as the old days (i.e. AS55_I, AS55_Q and REFL_I).(#4760)

But remember : this is just a beginning of several measurements and tests to characterize the central part.

 

Here is a list of the measurements and actions :

  - 3f locking related

      + Listing up the necessary RFPDs and their installations.

      + Calibration of the SRM actuator  => this is necessary to convert the sensing matrix into unit of [counts/m] or [W/m].

      + Measurement of the sensing matrix => check the performance of 3f signals. Also diagonalization of the LSC sensing matrix

      + Diagonalization of the output matrix.

      + Noise characterization of 3f PDs => confirm the noise are low enough to keep the lock of the central part

 - Power-recycling gain issue related (#5541)

     + Estimation of the mode matching efficiency => maybe we can use power-recycled ITMs to estimate it (?)

     + Implementation of auto alignment servos and scripts for MICH, PRCL and SRCL. => integrate it to the existing ASS model

     + Search for a possible loss factor

  5626   Thu Oct 6 15:40:57 2011 JenneUpdateLSCArm absl length data taken

[Katrin, Jenne]

We took the data for the new absolute length measurement of both arms, after the latest vent and move.  We will analyze soonly.  We had done a round of analysis,  but then Koji pointed out that our data wasn't so clean because the whitening filters were on (and saturated the ADC).  We now have the data (but not the analysis) for the better data with the WF off.

So our dirty-data preliminary number for the X arm is 37.73meters, which is 14cm different from our old length.  We were supposed to move by ~20cm, so....either this measurement is bad because the data sucked (which it did), or we are 6cm off.  Or both.

I'll do another analysis with the clean data for both arms later today/tomorrow.

  5625   Thu Oct 6 15:37:26 2011 JenneUpdateGreen LockingY-green Mech Shutter Button

[Katrin, Jenne]

We were poking around and tried to make a button for the Y-green shutter, just like the X-green already has.  I don't know where the X-green shutter button goes to in model-land, so I can't figure out if there is already a channel set up for the Y end.  Just switching the X for a Y didn't work.  Someone (maybe me) should fix this in the next soon.

  5624   Thu Oct 6 05:18:20 2011 kiwamuUpdateLSCNoise in AS55 was from clipping : fixed

It turned out the noise in AS55 was due to a clipping.  After fixing the clipping the noise successfully went down.

I was going to briefly check the clipping and go ahead locking DRMI, but for some reason I couldn't stop myself from working on this issue.

Here is a plot of the noise spectra taken before and after fixing the clipping.

The configuration of this measurement is exactly the same as that I did before (#5595)

outofloop.png

(what I did)

 + Locked power-recycled ITMY so that the AS beam is bright enough to work with.

 + Shook BS at 1 Hz in the YAW direction

 + Looked around the AP table with an IR viewer and searched for a clipping moving at 1Hz.

 + Found the first lens in the AS beam path has clipped the beam at the upper side. A tiny portion of the beam was clipped.

 + Corrected the beam height to 4 inch by steering the very first mirror.

 + Raised the height of the lens because it was about 3.5 inch or so.

 + Found the lens had a scratch (~1 mm size ) at 1 cm blow the center on the surface.

   => I tried finding a spare 2 inch lens with a long focul length, but I couldn't find it,

        So I left the lens as it is, but we should buy some 2 inch lenses just in case like this.

 + Replaced the 1 inch beam splitter by 2-inch 99% BS so that most of the light goes into the RFPD and a little bit goes into the camera.

Quote from #5595

The AS55 signal contains more noise than the REFL signals.

  5623   Wed Oct 5 18:31:02 2011 KatrinUpdateGreen LockingExchanged mirror on YARM table

On the Green YARM end table the second mirror behind the laser has been exchanged.

Reason: The light is impinging on the mirror under an angle of  about 10 degrees, but the old mirror was coated for angle of incidence (aoi) of 45°.

Thus it was more like a beam splitter. The new mirror is a 1" Goock & Housego mirror which has a better performance for an aoi of 10 degree.

Realignment through Faraday Isolator and SHG cristall as well as 532nm isolator is almost finished.

  5622   Wed Oct 5 17:08:49 2011 steveUpdateSUSBS oplev spectra

Kiwamu and Steve,

The He/Ne oplev shows no coherece so relative intensity noise is not limiting factor for the oplev servo

Attachment 1: BSoplservON2.png
BSoplservON2.png
  5621   Wed Oct 5 14:18:09 2011 kiwamuUpdateCDSsome DAQ channel lost in c1sus : fb, c1sus and c1pem restarted

I found again the ini files had been refreshed.

I ran the activateDQ.py script (link to the script wiki page) and restarted the daqd process on fb.

The activateDQ.py script should be included into the recompile or rebuild scripts so that we don't have to run the script everytime by hands.

I am going to add this topic on the CDS todo list (wiki page).

Quote from #5561

Somehow some DAQ channels for C1SUS have disappeared from the DAQ channel list.

 

  5620   Wed Oct 5 11:33:25 2011 steveUpdateSUSPRM and BS oplev laser replaced

 

JDSU 1103P died after 4 years of service. It was replaced with new identical head of 2.9 mW output. The power supply was also changed.

The return spots of 0.04 mW  2.5 mm diameter on qpds are BS  3,700 counts and PRM 4,250 counts.

 

  5619   Tue Oct 4 20:34:20 2011 KatrinUpdateGreen Locking7kHz Peak in servo input YARM

[Kiwamu, Katrin]

As reported earlier an oscillation around 7kHz is an the PDH error signal. The lower spectrum show that there is a peak from 6-7kHz.

This peak is somehow dependent on the modulation frequency. This means the peak can be shifted to a higher frequency when the modulation frequency is increased (see for comparsion f_mod=279kHz).

If the power supply for the green PD is switched of the peak vanishes. The same happens if the LO is switched of.

servoinput.png servoinput2.png

  5618   Tue Oct 4 19:31:17 2011 kiwamuUpdateSUSPRM and BS oplev laser died

The He-Ne laser which has been used for the PRM and BS oplevs were found to be dead.

According to the trend data shown below, it became dead during the dolphin issue.

(During the dolphin issue the output from the oplev QPDs are digitally zero)

oplevs.png

  5617   Tue Oct 4 19:06:46 2011 KojiUpdateIOOclosed the shutter before the MC

Finished the work at 6:30

MC REFL (INLOCK): 0.50-0.52
MC REFL (UNLOCK): 6.9
MC TRANS: 54400~547000


RFAM level

Before the work: -48.5dBm for 1.07VDC (both 50Ohm terminated)

Right after the work: -80dBm for 0.896VDC (both 50Ohm terminated)
10min after:   -70dBm
1hour after:   -65dBm
3hours after: -62dBm
1day after (Oct 5, 20:00):    -62.5dBm
2days after (Oct 6, 23:20): -72.5dBm
3 days after (Oct 7, 21:00): -57.8dBm

Quote:

The shutter before the MC was closed at 3:30 as I started working on the RFAM.

MC REFL (INLOCK): 0.6~0.7
MC REFL (UNLOCK): 6.9
MC TRANS: 50000~52000

 

  5616   Tue Oct 4 16:58:45 2011 SureshUpdatePSLAM / PM ratio

Correction: Koji noted that Mirko actually reports a PM modulation index of 0.17 for the 11 MHz sideband (elog: http://nodus.ligo.caltech.edu:8080/40m/5462. This means

f) the amplitude ratio of the PM side-band to carrier is half of that = 0.084

g)  the PM to AM amplitude ratio as 0.084 / [4.0 x 10^(-4)]  = 209.

  5615   Tue Oct 4 16:10:45 2011 SureshUpdateComputer Scripts / ProgramsMC was not damping

The MC was not damping earlier this morning around 11:45AM.  The reason was that the INMATRIX on all the three MC1, 2 and 3 were zero.   This was seen earlier when autoburt did not restore this.

This got fixed when I Burt-Restored c1mcsepics.snap

In the afternoon we had to restore c1mcs again to restore the MC_TRANS channels because its INMATRIX was also zero.

Can we do something to make sure this gets done in the autoburt properly?

  5614   Tue Oct 4 15:57:30 2011 SureshUpdateIOOMC Trans channels are digital 0

I thought this problem might be arising because the MC2_TRANS QPD signals are not being passed from the c1mcs to c1ioo models over the rfm.   But there was no way to check if there is any data being picked up in c1mcs model.  So I copied the MCTRANS block from the c1ioo model into the c1mcs.  This block takes the four segments of the MC2_TRANS QPD and computes the pitch, yaw and sum signals from that.   It also exports these into epics channels.  I then recompiled and started the c1ioo c1mcs and c1rfm models. 

Restared fb at

Tue Oct  4 15:19:10 PDT 2011

Koji then noted that the MC2_TRANS filter banks in c1rfm and in c1ioo were showing nonzero values.   So the signals were infact reaching the c1ioo model.  They were being blocked by the INMATRIX (which the autoburt had not restored) of the MC_TRANS block, because all its elements were zero.  We burtrestored the c1iooepics to about 30hrs ago and then MC_TRANS signals were back in the LOCK_MC screen.

 

  5613   Tue Oct 4 15:43:10 2011 KojiUpdateIOOclosed the shutter before the MC

The shutter before the MC was closed at 3:30 as I started working on the RFAM.

MC REFL (INLOCK): 0.6~0.7
MC REFL (UNLOCK): 6.9
MC TRANS: 50000~52000

  5612   Tue Oct 4 14:41:47 2011 JenneUpdateIOOMC Trans channels are digital 0

I relocked the PMC (why is it unlocking so much lately??), and then noticed that even though the MC is locked, MC Trans Sum, P, Y, are all seeing digital zero.  I'm putting Suresh, as IOO guy, in charge of figuring it out.

  5611   Tue Oct 4 10:35:08 2011 MirkoUpdateCDSc1lsc and c1sus running again

[Alex, Mirko]

Alex fixed the computers this morning. It was in fact a dolphin problem:

Hi Jenne,  figured it out. Even though dxadmin said the Dolphin net was fine, it wasn't. Something happeneed to DIS networkmanager and I had to restart it. It is running on fb: 
controls@fb ~ $ ps -e | grep dis 12280 ?        00:00:00 dis_networkmgr
controls@fb ~ $ sudo /etc/init.d/dis_networkmgr restart
Once the restart was done both c1lsc and c1sus nodes were configured correctly, they printed the usual "node 12 is OK" "node 8 is OK" messages into the dmesg and was able to run /etc/start_fes.sh on lsc and sus to load all the FEs.  Alex

Some lights on c1lsc were still red: C1:DAQ-FBO_C1SYS_SYS and the smaller red light left of it. Restarted the fb. Didn't help. Restarted c1lsc, all green now.
Restored autoburt from Oct 3. 19:07 on c1lsc and c1sus.
  5610   Tue Oct 4 07:51:36 2011 steveUpdateCDSc1lsc and c1sus are still down

 

 c1lsc and c1sus are still down. Only ETMX and ETMY are damped

  5609   Mon Oct 3 23:52:49 2011 kiwamuUpdateCDSsome more tests for the Dolphin

[Koji / Kiwamu]

 We did several tests to figure out what could be a source of the computer issue.

The Dolphin switch box looks suspicious, but not 100% sure.

 

(what we did)

 + Removed the pciRfm sentence from the c1x04 model to disable the Dolphin connection in the software.

 + Found no difference in the Makefile, which is supposed to comment out the Dolphin connection sentences.

   ==> So we had to edit the Makefile by ourselves

 + Did a hand-comilpe by editing the Makefile and running a make command.

 + Restarted the c1x04 process and it ran wihtout problems.

   ==> the Dolphin connection was somehow preventing the c1x04 process from runnning.

 +  Unplugged the Dolphin cables on the back side of the Dolphin box and re-plug them to other ports.

   ==> didn't improve the issue.

 + During those tests, c1lsc frequently became frozen. We disabled the automatic-start of c1lsc, c1ass, c1oaf by editting rtsystab.

  ==> after the test we reverted it.

 + We reverted all the things to the previous configuration.

  5608   Mon Oct 3 21:20:30 2011 JenneUpdateCDSc1lsc and c1sus didn't run

[Jenne, Mirko, Kiwamu, Koji, and Jamie by phone]

We just got off the phone with Jamie.  In addition to all the stuff that Kiwamu mentioned, Mirko reverted the c1oaf model and C-code to stuff that was working successfully on Friday (using "svn export, rev # 1134) since that's what we were working on when all hell broke loose.

We did a few rounds of "sudo shutdown -h now" on c1lsc and c1sus machines, and pulled the power cords out.  

We also switched the c1ioo and c1lsc 1PPS fibers in the fanout chassis, to see if that would fix the problem.  Nope.  c1ioo is still fine, and c1lsc is still not fine.

Still getting "No Sync".

We're going to call in Alex in the morning, if we can't figure it out soon.

  5607   Mon Oct 3 20:47:51 2011 kiwamuUpdateCDSc1lsc and c1sus didn't run

[Mirko / Jenne / Kiwamu]

Just a quick update. All the realtime processes on the c1lsc and c1sus machine didn't run at all.

Somehow the c1xxxfe.ko kernel module, where xxx is x04, x02, lsc, ass, sus, mcs, pem and rfm failed to be insmod.

The timing indicators on the c1lsc and c1sus machine are saying NO SYNC.

 

- According to log files (target/c1lsc/logs/log.txt)

insmod: error inserting '/opt/rtcds/caltech/c1/target/c1lsc/bin/c1lscfe.ko': -1 Unknown symbol in module

- dmesg on c1lsc (c1sus also dumps the same error message):

[   45.831507] DXH Adapter 0 : sci_map_segment - Failed to map segment - error=0x40000d01
[   45.833006] c1x04: DIS segment mapping status 1073745153

DXH dapter is a part of the Dolphine connections.

When a realtime codes is waking up, the code checks the Dolphin connections.

The checking procedure is defined by dolphin.c (/src/fe/doplhin.c).

According to a printk sentence in dolphin.c the second error message listed above will return status "0" if everything is fine.

The first error above is an error vector from a special dolphin's function called sci_map_segment, which is called in dolphin.c.

So something failed in this sci_map_segment function and is preventing the realtime code from waking up.

Note that sci_map_segment is defined in genif.h and genif.c which reside in /opt/srcdis/src/IRM_DSX/drv/src.

  5606   Mon Oct 3 20:02:59 2011 SureshUpdatePSLAM / PM ratio

[Koji, Suresh]

In the previous measurement, the PDA 255 had most probably saturated at DC, since the maximum ouput voltage of PDA255 is 5V when it is driving a 50 Ohm load.  It has a bandwidth of 0 to 50MHz and so can be reliably used to measure only the 11 MHz AM peak.  In this band it has a conversion efficiency of 7000 V per Watt (optical power at 1064nm).  [Conversion efficiency:  From the data sheet we get 0.7 A/W of photo-current at 1064nm and 10^4 V/A of transimpedance]  The transimpedance at 55 MHz is not given in the data sheet.  Even if PDA255 is driving a high impedance load, at high incident power levels the bandwidth will be reduced due to finite gain x bandwidth product of the opamps involved, so the conversion efficiency at 11 MHz would not be equal to that at DC.

So Koji repeated the measurement with a lower incident light level:

**********************************

V_DC = 1.07 V  with 50 Ohm termination on the multimeter.

Peak height at 11 MHz on the spectrum analyzer (50 Ohm input termination) = -48.54 dBm

***********************************

Calculation: 

a) RF_Power at 11 MHz :  -48.45 dBm = 1.4 x 10^(-8) W

b) RF_Power = [(V_rms)^2] / 50_ohm  ==> V_rms = 8.4 x 10^(-4) V

c) Optical Power at 11 MHz: [V_rms / 7000] = 1.2 x 10^(-7) W

d) Optical Power at DC =  [V_DC / 7000] = 1.46 x 10^(-4) W

e) Intensity ratio:  I_AM / I_c = 7.9 x 10^(-4) . AM:Carrier amplitude ratio is half of the intensity ratio = 4.0 x 10^(-4)

f) PM amplitude ratio from Mirko's measurement is 0.2

g) The PM to AM amplitude ratio is 506

_________________________________

As the AM peak is highly dependent upon the drifting EOM position in yaw, it is quite likely that a higher PM/AM ratio could occur.  But this measurement shows how small it could get if the current situation is allowed to continue.

 

Quote:

[Mirko / Kiwamu]

 We have reviewed the AM issue and confirmed the ratio of AM vs. PM had been about 6 x103.

The ratio sounds reasonably big, but in reality we still have some amount of offsets in the LSC demod signals.

Next week, Mirko will estimate the effect from a mismatch in the MC absolute length and the modulation frequency.

 


(Details)

 Please correct us if something is wrong in the calculations.

 According to the measurement done by Keiko (#5502):

        DC = 5.2 V

        AM @ 11 and 55 MHz = - 56 dBm = 0.35 mV (in 50 Ohm system)

Therefore the intensity modulation is 0.35 mV / 5.2 V = 6.7 x 10-5

Since the AM index is half of the intensity modulation index, our AM index is now about 3.4 x 10-5

According to Mirko's OSA measurement, the PM index have been about 0.2.

As a result,  PM/AM = 6 x 103

Quote from #5502

Measured values;

* DC power = 5.2V which is assumed to be 0.74mW according to the PDA255 manual.

*AM_f1 and AM_f2 power = -55.9 dBm = 2.5 * 10^(-9) W.

 

 

  5605   Mon Oct 3 17:57:12 2011 kiwamuUpdateASCIPPOS fixed

The input matrix of IPPOS were fixed so that the horizaontal motion correctly shows up in X and the vertial is Y.

 

(what I did)

 + The data base file, QPD.db, were edited.

  QPD.db is a part of the c1isxaux slow machine and it determines the input matrix for deriving the X/Y signals from each quadrant element.

 + The previous input matrix was :

      X = (SEG1 + SEG4) - (SEG2 + SEG3)

      Y = (SEG1 + SEG2) - (SEG3 + SEG4)

 + The new matrix which I set is :

     X = (SEG1 + SEG2) - (SEG3 + SEG4)

     Y = (SEG1 + SEG4) - (SEG2 + SEG3)

    The new matrix is a just swap of the previous X and Y.

 + Then c1isxaux was rebooted by :

     telnet  c1iscaux

     reboot          

 + The I did the burt restore it to this morning.

Quote from #5574

The channels for IPPOS had been assigned in a wrong way.

 

ELOG V3.1.3-